OpenStudy (anonymous):

2(4)(8)(16)= 1024 How?

OpenStudy (anonymous):

Use a calculator

myininaya (myininaya):

2*4=8 8*8=64 64*16=64*6+64*10=384+640=1024

OpenStudy (anonymous):

Parentheses means multiplication.

OpenStudy (anonymous):

Ty

OpenStudy (anonymous):

you mean T times y? :)

OpenStudy (anonymous):

D: i'm not sure math teacher just threw some homework at me! but it looks like that so.... Yes :)